OpenStudy (smurfy14):

Find the equation of the line in slope-intercept form for the line through (3, -2) with slope m=4/5

OpenStudy (anonymous):

It is known that: \[y = \frac{4}{5}x + c\]and \[\frac{4}{5}(3) + c = -2\]so \[c = -\frac{22}{5}\]therefore \[y = \frac{4}{5}x - \frac{22}{5}\]Alternatively: \[y = \frac{1}{5}(4x - 22)\]

OpenStudy (smurfy14):

alright and if i wrote it with the 1/5 distributed out would that still count as point intercept form?

OpenStudy (anonymous):

y= mx+b ===> y=(4/5)x -2

OpenStudy (katrinakaif):

Yes you have to distribute it to get a finalaized equation.

OpenStudy (katrinakaif):

*finalized

OpenStudy (smurfy14):

ok thanks!

OpenStudy (katrinakaif):

Anytime =)